solve the recurrence relation Pn = Pn-1 + n with initial condition P1 = 2 using interation.

1
Expert's answer
2021-11-29T11:18:14-0500

"P(n)=P(n-1)+n"


"P(n-1)=P(n-2)+n-1"


"P(n)=P(n-1)+n=P(n-2)+n-1+n"



"P(n)=P(n-2)+2n-(0+1)"



"P(n-2)=P(n-3)+n-2"

"P(n)=P(n-1)+n=P(n-2)+2n-1"

"=P(n-3)+3n-3"


"P(n)=P(n-3)+3n-(0+1+2)"



"P(n-3)=P(n-4)+n-3"


"P(n)=P(n-1)+n=P(n-2)+2n-1"


"=P(n-3)+3n-3=P(n-4)+4n-6"

"P(n)=P(n-4)+4n-(0+1+2+3)"



"..."

"P(n)=P(n-k)+kn-\displaystyle\sum_{i=0}^{k-1}i, \ k=1,2,..., n-1"

"P(1)=P(n-(n-1))"

Then


"P(n)=P(1)+(n-1)n-\displaystyle\sum_{i=0}^{n-2}i, n=2, 3, ..."

"\displaystyle\sum_{i=0}^{n-2}i,=\dfrac{(n-2)(n-2+1)}{2}"

"=\dfrac{(n-2)(n-1)}{2}, n=2, 3, ..."


"P(n)=2+(n-1)n-\dfrac{(n-2)(n-1)}{2}"

"=2+\dfrac{(n-1)(2n-n+2)}{2}=2+\dfrac{(n-1)(n+2)}{2}"

"=2+\dfrac{(n-1)(n+2)}{2}, n\geq2"

"P(n)=2+\dfrac{(n-1)(n+2)}{2},n\geq1"